OpenStudy (anonymous):

The minimum value of \(\Large \frac{(x^4+1)(y^4+1)(z^4+1)}{xy^2z}\) as x,y and z range over the positive reals is equal to \(\Large \frac{A\sqrt{B}}{C}\), where A and C are coprime and B is squarefree. What is \(\Large A+B+C\ ?\)

OpenStudy (vishweshshrimali5):

We would have to use inequalities

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Any idea ?

OpenStudy (anonymous):

Using AM-GM

OpenStudy (vishweshshrimali5):

Have you tried that ?

OpenStudy (vishweshshrimali5):

Gotcha !

OpenStudy (vishweshshrimali5):

Just a minute

OpenStudy (vishweshshrimali5):

I can write the given expression as: \[\Large {(x^3+\cfrac{1}{x})(y^2+\cfrac{1}{y^2})(z^3+\cfrac{1}{z})}\]

OpenStudy (vishweshshrimali5):

Now I can apply AM - GM to the second one

OpenStudy (vishweshshrimali5):

For first and third, I would have to think something...

OpenStudy (anonymous):

\(\Large x^3+\frac{1}{x} =x^3+\frac{1}{3x}+\frac{1}{3x}+\frac{1}{3x}\) and using AM-GM

OpenStudy (vishweshshrimali5):

Yup perfect

OpenStudy (vishweshshrimali5):

If you use AM - GM to all three, then it would simplify to : \[\large{\cfrac{32\sqrt{3}}{9}}\]

OpenStudy (vishweshshrimali5):

Co prime - check square free - check

OpenStudy (vishweshshrimali5):

A+B+C = 44

OpenStudy (vishweshshrimali5):

Done !

OpenStudy (vishweshshrimali5):

Another method - calculus !

OpenStudy (vishweshshrimali5):

Have you studied calculus @JungHyunRan ?

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Great!

OpenStudy (vishweshshrimali5):

Then using partial differentiation and maxima - minima, you can come up with the same answer more easily

OpenStudy (anonymous):

^^ I see. Thank you so much @vishweshshrimali5

OpenStudy (vishweshshrimali5):

No problem and it was entertaining to solve such problems after a very long time :)

OpenStudy (vishweshshrimali5):

Remember such problems which involve symmetry (as in here), calculus provides the shortest answers

ganeshie8 (ganeshie8):

I had seen AM-GM abused this nicely a couple of days back... for the first time, by @satellite73 :)

OpenStudy (vishweshshrimali5):

:D @satellite73 ROCKS !!

OpenStudy (vishweshshrimali5):

He has not got a 100 smart score for nothing after all !

OpenStudy (vishweshshrimali5):

@JungHyunRan, if you like such type of problems I would suggest you to go through some olympiad level books. They abuse very very simple inequalities in interesting ways :)

OpenStudy (anonymous):

yup

OpenStudy (vishweshshrimali5):

A great collection of books can be found at this link: http://www.imomath.com/index.php?options=347&lmm=0

ganeshie8 (ganeshie8):

im still not able to get to that problem solved by satellite, my answered questions are not loading >.< will post it when i find it :)

OpenStudy (vishweshshrimali5):

Sure @ganeshie8 :)